OpenStudy (anonymous):

Intro to Statistics: \[\mu=\sum xP(x)\] \[\mu\] is the expected value of x So would I add the x values and P(x) values and then multiply them to get \[\mu\]? x: 0 1 2 3 4 5 P(x): 0.237 0.396 0.264 0.088 0.015 0.001

hartnn (hartnn):

multiply corresponding x and p(x) values, like 0*0.237+1*0.396 .......

OpenStudy (slaaibak):

you would multiply 0 by 0.237 add it to 1 times .396 add it to 2 times 0.264 etc... the sum would be u

OpenStudy (anonymous):

oh ok. It seemed too simple

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

for the standard deviation I have the equation \[\sigma=\sqrt{\sum(x-\mu)^2P(x)}\] So, here I do: \[0-\mu=-1.253\] \[1-\mu=1-1.253=-.253\] \[2-\mu=2-1.253=.747\] \[3-\mu=3-1.253=1.747\] \[4-\mu=4-1.253=2.747\] \[5-\mu=5-1.253=3.747\] square each solution then multiply to P(x), then add those solutions up

hartnn (hartnn):

yes! finally take square root then.

OpenStudy (anonymous):

oh yes of course. Thank you!

OpenStudy (anonymous):

how do I make a table in latex? I have to submit my work electronically and a table would make it easier.

OpenStudy (anonymous):

I couldn't find it here http://en.wikipedia.org/wiki/Help:Displaying_a_formula#Diagrams_in_TeX

hartnn (hartnn):

let me search.

hartnn (hartnn):

maybe write in matrix form .....of nX2 i'll look up something specifically for table

OpenStudy (anonymous):

oh ok, thanks

OpenStudy (anonymous):

\hlines what does that do?

OpenStudy (anonymous):

I think I get it. I'll just use \\ \hlines before i move on to the next line. Makes sense

hartnn (hartnn):

ok and http://en.wikibooks.org/wiki/LaTeX/Tables

OpenStudy (anonymous):

oh perfect! There's the explanations for \hlines. Thanks! =D